OpenStudy (anonymous):

find the missing angle measure in isoceles triangle XYZ, with vertext angle at Y. angle Y= 70 degrees, angle Z=

OpenStudy (jennychan12):

an isoceles has two equal angles. since the vertex angle is 70 degrees the other two angles must be the same and you know that a triangle has 180 degrees try it!

OpenStudy (anonymous):

I got 110 is that right?

OpenStudy (anonymous):

its (180-70)/2=110/2=55

OpenStudy (jennychan12):

110 is <X + <Z divide 110 in half

OpenStudy (anonymous):

thalli

OpenStudy (jennychan12):

it's 55 for <X and 55 for <Z